OpenStudy (anonymous):

Could someone explain this please? My calculus teacher wrote that ln(p) - ln (m-p) = kt +c leads to the step - ln ((m-p)/p) = kt + c. I thought the next step would be ln( p/(m-p) ).

OpenStudy (mertsj):

\[\ln p-\ln (m-p)=-1[-\ln p+\ln (m-p)]=-1[\ln \frac{m-p}{p}]\]

OpenStudy (anonymous):

Oh, I see! Thank you! :)

OpenStudy (mertsj):

yw